OpenStudy (anonymous):

Which of the following is equivalent to the expression (3ab)(-5ab)? A. -2ab B. -2a^2b^2 C. -15ab D. -15a^2b^2 When solving this problem i got A. but the answer is D. I want to know what i did wrong and how to solve this.

OpenStudy (anonymous):

D is the correct answer

OpenStudy (anonymous):

okay well let's see what your work was to get A

OpenStudy (anonymous):

3*-5 = -15 a*a = a^2 b*b = b^2

OpenStudy (anonymous):

you're multiplying them, not adding them

OpenStudy (anonymous):

(3ab)(-5ab) 3 * -5 = 15 and a * a = a^2 and b * b = b^2

OpenStudy (anonymous):

does our work help

OpenStudy (anonymous):

do you always multiply the exponents?

OpenStudy (anonymous):

no, you are adding the exponents a^1 * a^1 = a^2